OpenStudy (aaronandyson):

matrix 1 by 2 is ( x y) times matrix 2 by 2 is (-2 2 1 3) equals matrix (-3 5) find x and y

OpenStudy (aaronandyson):

@ganeshie8

OpenStudy (thomas5267):

\[ \begin{bmatrix} x&y \end{bmatrix} \begin{bmatrix} -2&2\\ 1&3 \end{bmatrix}= \begin{bmatrix} -3&5 \end{bmatrix}? \]